# On the mechanism of black hole energy reduction in the Blandford-Znajek process

### v1

@@ 5.1 Previous studies on the BH energy reduction
Toma and Takahara toma16 discussed the time-dependent behavior of the front of magnetically-dominated inflow, but they assumed a vacuum between the horizon and the inflow, instead of the membrane. The existence of the membrane is more natural than the vacuum. Although they provided with the key equations, Eqs. (8) and (10), at the front of the inflow, they insisted on seeking properties invariant between the BL and KS coordinates. They did not find the mechanism of BH energy reduction by formation of $e_{\rm in}<0$, because $e_{\rm in}$ is positive in the KS coordinates.
Infall of negative energy objects
The problem what kind of negative energy falls into the BH in the BZ process was first discussed in depth by Komissarov komissarov09 . The argument there tried to find a certain equation $S^{r}=ev^{r}_{\rm EM}$ for the entire inflow instead of only the front of inflow, but the inflow velocity $v^{r}_{\rm EM}$ was not identified. Takahashi et al. takahashi90 argued that $S^{r}\approx\mathcal{E}\alpha\rho u^{t}v^{r}$ (Eq. 23) means a negative energy infall with $\mathcal{E}<0$ and $v^{r}<0$ for the entire inflow, but $\mathcal{E}$ is related to $e$ only at the front of inflow. Koide and Baba koide14 wrote $S^{\hat{r}}=\alpha ev^{r}_{\rm EM}$ at the horizon similarly to Eq. (38), but they interpreted $v^{r}_{\rm EM}$ as the steady particle drift velocity measured by the FIDO, $D_{\hat{\theta}}B_{\hat{\varphi}}/B^{2}$. This interpretation is strange since $e$ is the quantity in the coordinate basis. Kinoshita and Igata kinoshita18 developed the formalism of world sheets of rigidly rotating Nambu-Goto strings, which has properties similar to the field sheets of forc-free electromagnetic fields, and suggested that the ‘stretching of the magnetic field lines’ plays the role of the infalling objects. This might correspond to the front of the MHD inflow in our picture. An analogy of the Alfvén wave superradiance to the BZ process has been also investigated noda22 . It is beyond the scope of this paper to proceed with the discussion in this direction.
@@ 5.2 Driving force of poloidal current
The remaining problem is what drives the poloidal electric current maintaining $H_{\varphi}$. In the unipolar induction of the pulsar winds, the $\mathbf{v}\times\mathbf{B}$ force in the matter-dominated star drives the poloidal current in the direction of $-\mathbf{E}$, which drives the steady polidal current circuit. In contrast, the BZ process has no matter-dominated region, and the cross-field current cannot flow effectively in the inflow or the outflow. The charged particles carrying the current are provided from the outside into the ordered magnetic field of the BH magnetosphere (e.g. levinson11, ; kimura22, ), which must be adjusted to carry the poloidal current and maintain $H_{\varphi}$.
Suppose that the conductive cross-field current was flowing on the membrane as in the membrane paradigm, then one would consider that the cross-field current in the opposite direction is driven somewhere, for example at the separation surface or the surface of $\Omega=\Omega_{\rm F}$ okamoto05 . However, we have shown that $\eta^{\theta}\approx 0$ on the falling membrane, and thus no cross-field current is required to be driven. The magnetically dominated inflow and outflow has no circuit of poloidal electric current, and the poloidal current just flows along the magnetic field lines.
@@ Acknowledgment
K. T. report with sadness that our coauthor Fumio Takahara, who contributed seminal ideas to this work, passed away 2023 June 6. K. T. thanks T. Harada, M. Kimura, A. Naruko, K. Ogasawara, S. J. Tanaka, S. Kisaka, and K. Kin for useful discussions. This work is partly supported by JSPS Grants-in-Aid for Scientific Research No. 18H01245 (K. T.) and No. 24K07100 (M. N.).

### v2

@@ 5.1.2 Different views in the BL and KS coordinates
Toma and Takahara toma16 discussed the time-dependent behavior of the front of magnetically-dominated inflow, but they assumed a vacuum between the horizon and the inflow, instead of the falling membrane. The existence of the falling membrane is more natural than the vacuum. Although they provided with the key equations, Eqs. (8) and (10), at the front of the inflow, they did not find the mechanism of BH energy reduction by formation of $e_{\rm in}<0$. This was partly because they insisted on seeking properties invariant between the BL and KS coordinates, and $e_{\rm in}$ is positive in the KS coordinates.
5.1.3 Infall of negative energy objects
The problem what kind of negative energy falls into the BH in the BZ process was first discussed in depth by Komissarov komissarov09 (see also lasota14, ). The argument there utilized the BL coordinates and tried to find a certain equation $S^{r}=ev^{r}_{\rm EM}$ for the part of the inflow with $D^{\theta}>0$ (i.e., $\Omega>\Omega_{\rm F}$) instead of only the front of the inflow, but the infall velocity $v^{r}_{\rm EM}$ was not identified. It was emphasized that the electromagnetic energy flux measured by the FIDOs, $S^{r}_{\rm FIDO}=D_{\theta}B_{\varphi}/4\pi\sqrt{\gamma}$, is in the opposite direction of $S^{r}=E_{\theta}H_{\varphi}/4\pi\sqrt{\gamma}$ where $\Omega>\Omega_{\rm F}$. However, $S^{r}_{\rm FIDO}$ is not straighforwardly related to the electromagnetic energy density in the coordinate basis $e$ (‘energy density measured at infinity’). In contrast, Eq. (36) is constituted of only quantities in the coordinate basis, $S^{r}_{\rm in},e_{\rm in},$ and $V$.
Koide and Baba koide14 wrote $S^{\hat{r}}=\alpha ev^{r}_{\rm EM}$ at the horizon similarly to Eq. (36), but they interpreted $v^{r}_{\rm EM}$ as the steady particle drift velocity measured by the FIDOs, $D_{\hat{\theta}}B_{\hat{\varphi}}/B^{2}$. This interpretation is similar to that in komissarov09 .
Takahashi et al. takahashi90 argued that $S^{r}\approx\mathcal{E}\alpha\rho u^{t}v^{r}$ (Eq. 20) means a negative energy infall with $\mathcal{E}<0$ and $v^{r}<0$ for the entire inflow. However, $\mathcal{E}$ is related to $e$ only at the front of inflow. Moreover, all the above interpretations in komissarov09 ; koide14 ; takahashi90 tried to find an equation of type $S^{r}=ev^{r}_{\rm EM}$ for the steady part of the inflow. However, $S^{r}$ is the product of the steady fields $E_{\theta}$ and $H_{\varphi}$, so that its velocity could not be defined. In our picture, we have considered the velocity $V$ of the inflow front, which is unsteady part of the inflow.
Kinoshita and Igata kinoshita18 developed the formalism of world sheets of rigidly rotating Nambu-Goto strings, which has properties similar to the field sheets of force-free electromagnetic fields, and suggested that the ‘stretching of the magnetic field lines’ plays the role of the infalling objects. This might correspond to the front of the MHD inflow in our picture.
An analogy of the Alfvén wave superradiance to the BZ process has also been investigated noda22 . It is beyond the scope of this paper to proceed with the discussion in this direction.
@@ 5.2 Driving force of poloidal current
The remaining problem is what drives the poloidal electric current maintaining $H_{\varphi}$. In the unipolar induction of the pulsar winds, the $\mathbf{v}\times\mathbf{B}$ force in the matter-dominated star drives the poloidal current in the direction of $-\mathbf{E}$, which drives the steady poloidal current circuit. In contrast, the BZ process has no matter-dominated region, and the cross-field current cannot flow effectively in the inflow or the outflow as seen in Section 3. The charged particles carrying the current are provided from the outside into the ordered magnetic field of the BH magnetosphere (e.g. levinson11, ; kimura22, ), which must be adjusted to carry the poloidal current and maintain $H_{\varphi}$.
Suppose that the conductive cross-field current (Eq. 37) was flowing on the membrane as in the membrane paradigm, then one would consider that the cross-field current in the opposite direction is driven somewhere, for example at the separation surface or the surface of $\Omega=\Omega_{\rm F}$ okamoto05 . However, we have shown that $\eta^{\theta}\approx 0$ on the falling membrane, and thus no cross-field current is required to be driven. The magnetically dominated inflow and outflow has no circuit of poloidal electric current, and the poloidal current just flows along the magnetic field lines.
@@ Acknowledgment
K. T. reports with sadness that our coauthor Fumio Takahara, who contributed seminal ideas to this work, passed away 2023 June 6. K. T. thanks T. Harada, M. Kimura, A. Naruko, K. Ogasawara, S. J. Tanaka, S. Kisaka, and K. Kin for useful discussions. This work is partly supported by JSPS Grants-in-Aid for Scientific Research No. 18H01245 (K. T.) and No. 24K07100 (M. N.).
@@
These equations include $\eta$ only as $H_{\varphi}/\eta$ and $B_{\rm p}/\eta$ since $M^{2}=4\pi\rho\alpha^{2}u_{\rm p}^{2}/B_{\rm p}^{2}=\alpha u_{\rm p}/(B_{\rm p}/4\pi\eta)$. Therefore $\eta$ does not affect the structure of flow solution, but just affects the normalizations of field strengths.
Appendix B MHD inflow Solutions for $\theta=\pi/6$
Figure 4: Characteristic radii of our MHD inflow solutions for $\theta=\pi/6$. We plot the radii of separation point (thick solid line), inner Alfvén point (square), inner magnetosonic point (black dot), point of $e=0$ (thin solid line), that of $|v^{r}|/(\alpha/\sqrt{\gamma_{rr}})=0.99$ (dashed line), and that of $D_{\hat{\theta}}/|B_{\hat{\varphi}}|=0.99$ (dot-dashed line) for the case of each $\mathcal{E}$. The parameter sets for the top left, top right, bottom left, and bottm right panels are $(a,\Omega_{\rm F}/\Omega_{\rm H},\mathcal{E}-\mathcal{L}\Omega_{\rm F})=(0.8,0.5,0.55)$, $(0.5,0.5,0.55)$, $(0.8,0.4,0.55)$, and $(0.8,0.5,0.45)$, respectively.
For reference, we show the characteristic radii of the steady, axisymmetric, cold MHD inflow solutions along a split-monopole poloidal magnetic field line with $\theta=\pi/6$ in Figure 4. For $\theta=\pi/6$, the ergosphere is inside the radius of $r_{\rm es}-r_{\rm H}\approx 0.15$ for $a=0.8$ and $r_{\rm es}\approx 0.03$ for $a=0.5$.